Consider the following two bonds. One bond with a coupon rate of 4%, semi-annual coupons, and...

  1. Consider the following two bonds. One bond with a coupon rate of 4%, semi-annual coupons, and 10 years until maturity. The second bond has 5 years until maturity but is otherwise the same.
    1. What is the most you should pay for each asset if current yields are 6%?
    2. Do the bonds sell at a premium or a discount?
    3. Suppose current yields increase to 7%, what are the new bond prices?
    4. Which bond is more sensitive to yield changes? Why?
Answer #1

Number of periods to maturity = n

Yield to Maturity = r

Semi Annual Coupon Payment P

Face Value = FV

Hence, Present Value of bond = PV = P/(1+r) + P/(1+r)2 + .... + P/(1+r)n + FV/(1+r)n

= P[1 - (1+r)-n]/r + FV/(1+r)n

(a)

Bond 1
n = 10*2 = 20 semiannual periods
r = 0.06/2 = 0.03
FV = $100
P = 4%*100/2 = $2
Hence, PV = 2[1 - (1+0.03)-20]/0.03 + 100/(1+0.03)20 = $85.12

Bond 2
n = 5*2 = 10 semiannual periods
r = 0.06/2 = 0.03
FV = $100
P = 4%*100/2 = $2
Hence, PV = 2[1 - (1+0.03)-10]/0.03 + 100/(1+0.03)10 = $91.47

(b) Since the present value of both the bonds is less than the par value (FV) of the bond, both sell at discount

(c)

Bond 1
n = 10*2 = 20 semiannual periods
r = 0.07/2 = 0.035
FV = $100
P = 4%*100/2 = $2
Hence, PV = 2[1 - (1+0.035)-20]/0.035 + 100/(1+0.035)20 = $78.68

Bond 2
n = 5*2 = 10 semiannual periods
r = 0.07/2 = 0.035
FV = $100
P = 4%*100/2 = $2
Hence, PV = 2[1 - (1+0.03)-10]/0.03 + 100/(1+0.03)10 = $87.53

(d) The Bond with 10 years to maturity is more sensitive to yield changes. The longer the maturity of a bond, the more sensitive it is to yield changes.